The most recent financial statements for Heine, Inc., are shown here: Income Statement Sales Costs Balance Sheet $28,400 Assets 59,300 Debt $25,500 33,800 19,500 Equity Taxable income $ 8,900 Tota$59,300 $59,300 Total $59,300 Taxes (40%) 3,560 Net income 5,340 Assets and costs are proportional to sales. Debt and equity are not. A dividend of $2,400 was paid, and the company wishes to maintain a constant payout ratio. Next years sales are projected to be $32,660. What is the external financing needed? (Do not round intermediate calculations.)
Answer #1

Growth rate in sales=(32660-28400)/28400=15%

Dividend payout ratio=Dividend/Net income

=(2400/5340)=0.449438202

Sales 32660
Costs(19500*1.15) $22425
Taxable income $10235
Taxes(40%)($10235*40%) $4094
Net income $6141
Less:Dividends($6141*0.449438202) $2760
Addition to retained earnings $3381

Total assets would be=$59300*1.15=$68195

Total equity=$33800+Addition to retained earnings

=$33800+3381=$37181

Total assets=Total equity+Total liabilities

Hence external financing needed=$68195-$37181-$25500

=$5514.